OpenStudy (anonymous):

Water has an index of refraction of 1.33. What is the critical angle for light leaving a pool of water into air? 0°, 90°, 37° , or 49° ? :/

OpenStudy (michele_laino):

I remember that the critical angle is given by the subsequent formula: \[\Large \sin \left( {{\theta _c}} \right) = \frac{1}{n} = \frac{1}{{1.33}}\]

OpenStudy (anonymous):

okay! so that gets 0.751879 right? what happens next?

OpenStudy (michele_laino):

we have to compute this value: \[\Large \arcsin \left( {0.7518} \right) = ...\]

OpenStudy (anonymous):

0.850787638?

OpenStudy (michele_laino):

no, it is: 48.753

OpenStudy (anonymous):

ohh :/ oops :P okay! so that means our solution is 49º?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! thank you!

OpenStudy (michele_laino):

:)